OpenStudy (anonymous):

in a certain game, a player throws a standard 6 sided number cube twice per turn. what is the probability that the results of the first throw will be a 5 or a 6, and the second throw will be an even number?

OpenStudy (anonymous):

3214266589

OpenStudy (anonymous):

Timothy Hinds Sea Shell drive, merritt island florida 32952 social security 350-78-9989

OpenStudy (anonymous):

He enjoys long walk on his mothers tits and loves to cross dress and worship satan

OpenStudy (anonymous):

Since both events are independent, we have that, the sum of the events over the total number of events gives the probability (for each individual one). Hence, the probability of this happening is: \[ \left(2\cdot \frac{1}{6}\right)\cdot\left(3\cdot \frac{1}{6}\right)=\frac{1}{6} \]Since 5 and 6 are two of six possibilities and an even number is three of six.